A sample of cells has a total receptor concentration of 24 μM, and 35% of the receptors are occupied with ligand. The concentration of free ligand is 10 μM. Calculate Kd. { Kd = [R][L]/[R·L]}.
Answer #1

The total receptor concentration {R} is 24 micromolar

Total free ligand are =10 Micromolar

35% of the receptor are occupied with ligand

It means 65% of the receptor is 10

1% of the receptor is 10/65= 0.1538

100% of the receptor {L} is =0.1538×100= 15.38 micromolar.

Now the concentration of 35% of the receptor is

(35/100) × 15.38=5.38 micromolar

Now kd ={ [R] [L]/[R.L]}

Kd= {24×15.38/5.38}

Kd= 368.4/5.38=68.85

Kd = 68.85